# Reconstructing KV Caches with Cross-Layer Fusion for Enhanced Transformers

### v1

@@ 1 Introduction
Large language models (LLMs) based on the Transformer (vaswani2017attention) decoder have achieved unprecedented success across a wide range of tasks and applications. However, practical deployment for real-world applications requires LLM inference to achieve low latency and high throughput (hu2024minicpm; ge2024little; dao2022flashattention). This becomes particularly challenging with increasingly long contexts, primarily due to the linear memory overhead of the key-value (KV) caches in autoregressive generation (liu2024minicache; adnan2024keyformer; li2024survey).
To mitigate this, various methods have been proposed to reduce KV caches. Within the layer, Group-Query Attention (GQA  [^2] Unless otherwise noted, GQA mentioned in this paper uses a configuration where two query heads share a single key-value head.(ainslie2023gqa) and Multi-Query Attention (MQA) (shazeer2019fasttransformerdecodingwritehead) reduce redundancy by sharing one head’s key and value across multiple query heads. MLA (liu2024deepseek) uses low-rank joint compression for key-value caches. Across the layer, CLA (brandon2405reducing) shares the KV cache between adjacent layers, while YOCO  [^3] Note that in the original YOCO paper, the cache-storing layers use efficient attention mechanisms such as Sliding Window Attention(SWA). In this paper, unless otherwise stated, we use full attention. (sun2024yoco) reuses the intermediate layer cache for the top-half layers. While cross-layer sharing techniques also reduce memory overhead, they consistently underperform within-layer methods, which motivates the need for a more effective cross-layer sharing strategy.
@@ 2.2 Reconstruction Function $\mathcal{F}$
the reconstruction function $\mathcal{F}_{i}$ acts as a selector, the source layer mapping $\Phi(i)$ typically contains a single index, i.e., $|\Phi(i)|=1$. Recent methods such as YOCO (sun2024yoco) and CLA (brandon2405reducing) adopt this approach.
@@ 2.4 RoPE Compatibility
To ensure that learnable weights do not disrupt the relative positional encoding property of RoPE (su2024roformer), we analyze the attention score. For a query at position $m$ and a key at position $n$, the attention score $A_{n,j}$ in the $j$-th 2D subspace, after incorporating a learnable weight vector ${\bm{w}}_{j}=[w_{2j},w_{2j+1}]^{T}$, can be decomposed as follows (see Appendix A.4 for the full derivation):
@@
Method Prefilling FLOPs Decoding FLOPs
Cache
Mem.
Cache
I/O.
@@ 3 Experiments
We introduce three dense language models with 332M, 650M and 1.5B parameters, both following the Qwen3 architecture (qwen3technicalreport). More configuration are detailed in Table 6. All three models share the same configuration: a vocabulary of 128,000 tokens, a context length of 8,192, and 16 attention heads. They were trained on the FineWeb-Edu dataset (lozhkov2024fineweb-edu). We use the AdamW optimizer (loshchilov2017decoupled) with $\beta_{1}=0.9$, $\beta_{2}=0.95$. The learning rate followed a cosine schedule, warming up for 2,000 steps to a peak of $3\times 10^{-4}$ and decaying to a minimum of $3\times 10^{-5}$. A more detailed configuration can be found in Table 4.
@@
For model evaluation and comparison, we assess perplexity on a 500M-token validation split randomly sampled from FineWeb-Edu. We also evaluate perplexity on WikiText (merity2016pointer). We also conduct five-shot evaluation across diverse language tasks, including MNLI (williams2018broadcoveragechallengecorpussentence), SCIQ (welbl2017crowdsourcing), LAMBADA (LAMB-Acc) (paperno2016lambada), Hellaswag (zellers2019hellaswag), ARC-Easy (clark2018think), ARC-Challenge (clark2018think) and MMLU (hendryckstest2021). All evaluations are based on the LM Evaluation Harness (gao12608602framework).
@@
Model
Cache
Mem. $\downarrow$
Valid
Loss $\downarrow$
Wiki
Text $\downarrow$
MNLI
SCIQ
LAMB-
Acc
Hella
Swag
ARC-E
ARC-C
MMLU
Avg
@@ 3.2 Main Results
Model
Valid
Loss $\downarrow$
Wiki
Text $\downarrow$
MNLI
SCIQ
LAMB-
Acc
Hella
Swag
ARC-E
ARC-C
MMLU
Avg
@@ 5.1 Cross-Layer KV Cache
KV caching is a major memory bottleneck in LLM inference (han2025polarquantquantizingkvcaches; liao2024SA). Memory footprint can be reduced by limiting KV state retention to a subset of layers. YOCO (sun2024yoco) reuses a global KV cache from an intermediate layer for the model’s latter half, roughly halving memory; CLA (brandon2405reducing) shares caches across adjacent layers. More aggressive cross-layer reuse includes Mixture of Recurrent (bae2025mor), which stacks recurrent blocks (each comprising multiple transformer layers) so later blocks reuse the first block’s KV. SVFormer (SVFormer25), which does not share the key but reuses the bottom layer’s values; and LCKV (wu2024layer), which caches only the last layer’s KV for all earlier layers, with safeguards against cyclic dependencies. However, most approaches often overlook the distinct roles of keys and values: keys drive relevance scoring, while values supply contextual content. Treating KV as a single unit for caching neglects this functional disparity. Moreover, sharing KV states indiscriminately across layers risks undermining each layer’s specific role in hierarchical processing, potentially introducing irrelevant context and hindering its unique function. These critical aspects remain underexplored.
@@ 5.2 Early-Feature Fusion
Another line of research leverages early-layer feature fusion to stabilize gradients and enrich global features( (he2016deep; huang2017densely; pagliardini2024denseformer; abdullaev2025transformer)). ResNet (he2016deep) introduces residual connections that add outputs from previous layers to the current layer to stabilize training. Subsequent approaches further enhance the information flow within this framework. DenseNet (huang2017densely) concatenates the outputs of all previous layers, while Denseformer (pagliardini2024denseformer) and MambaMixer (behrouz2024mambamixer) incorporate learnable weighted coefficients into the residual paths, making the residual representations more expressive. Furthermore, reusing features from earlier layers can help alleviate the over-smoothing problem (wang2022anti; shi2022revisiting), where representations converge and lose diversity as the depth of the network increases, leading to degraded performance. To mitigate this issue, NeuTRENO (nguyen2023NeuTRENO) introduces a skip connection that injects a fraction of the bottom-layer output into every subsequent self-attention layer to mitigate over-smoothing. Similarly, (abdullaev2025transformer) enriches the transformer’s representational capacity by reusing value-residual information through a twice-attention mechanism.
@@ A.2 Architectures of Different Models
we adopt a decoder-only transformer architecture akin to Qwen3 (qwen3technicalreport), with model sizes ranging from 332M to 4B parameters. The specific architecture of models is summarized in Table 6. The hyperparameters for Pretraining is shown in Table 4.
@@
method
Valid
Loss $\downarrow$
Wiki
Text $\downarrow$
MNLI
SCIQ
LAMB-
Acc
Hella
Swag
ARC-E
ARC-C
MMLU
Avg
@@ A.6.1 Compatibility with MLA
While our proposed cross-layer fusion method, FusedKV, offers a novel way to reduce KV cache size by reconstructing information, alternative approaches achieve this through fundamental architectural modifications. A prominent example is Multi-Head Latent Attention (liu2024deepseek) , which reduces cache dimensionality by design rather than relying on reconstruction across layers. To investigate whether these two architextures are synergistic, we evaluate the compatibility and combined benefits of applying our method on top of an MLA-based model.
@@
method
Cache
Mem.
Wiki
Text $\downarrow$
MNLI
SCIQ
LAMB-
Acc
Hella
Swag
ARC-E
ARC-C
MMLU
Avg
@@
A direct empirical measurement of inference speed is heavily dependent on highly optimized, hardware-specific kernels. To provide a principled and timely analysis without the substantial engineering effort required for integration into production-grade frameworks like vLLM or SGLang, we estimated the inference speed using the open-source InferSim toolkit InferSim. We analyze two key metrics: Time to First Token (TTFT) for the prefill phase and Time Per Output Token (TPOT) for the decoding phase.
@@
Input/Output Length
TTFT (ms)
(MLA / MLA+FusedKV / MHA+FusedKV)
TPOT (ms)
(MLA / MLA+FusedKV / MHA+FusedKV)
@@
As established in recent work wang2025step; chang2024flux, when a model’s Arithmetic Intensity(AI) exceeds the GPU’s compute-to-memory ratio, the latency from I/O communication can be effectively hidden by computation. For MLA, the AI is approximately $4\times\text{heads}$ wang2025step. To demonstrate this principle, we configured a model with 48 attention heads, creating a compute-bound scenario where the AI of MLA+FusedKV surpasses NVIDIA H20 GPU’s compute-to-memory ratio. As shown in Table 10, in this setting, the additional I/O latency from our fusion step is fully overlapped by the increased computation. Consequently, while the prefill (TTFT) benefits remain, the decoding speed (TPOT) becomes identical to that of the standard MLA model.
@@ A.6.2 Compatibility with GQA
Beyond MLA, another widely adopted architectural optimization for improving inference efficiency is Grouped-Query Attention (GQA) ainslie2023gqa. To investigate synergistic effects with GQA, we conduct experiments on a 650M-parameter model, combining FusedKV with a GQA architecture. All training parameters were aligned with those used in Table 2. Our results, presented in Table 11, demonstrate that FusedKV is highly compatible with GQA and their combination yields a compelling performance-cost trade-off.
@@
method
Wiki
Text $\downarrow$
MNLI
SCIQ
LAMB-
Acc
Hella
Swag
ARC-E
ARC-C
MMLU
Avg
@@ A.6.4 Compatibility with sliding window
A prominent hybrid architecture is the integration of Sliding Window Attention (SWA) jiang2023mistral7b with full attention. In this section, we investigate the adaptability of FusedKV and FusedKV-Lite to a hybrid model that combines both SWA and full attention.
@@ A.6.5 COMPATIBILITY WITH OTHER EFFICIENCY TECHNIQUES
A key measure of a new efficiency technique’s utility is its ability to compose with existing methods. To evaluate the broader applicability of FusedKV and FusedKV-Lite, we investigate its compatibility with two prevalent model compression strategies: quantization (liu2024kivi) and pruning (sunsimple).
@@
Quantization reduces memory footprint by representing weights and activations with lower-precision data types. We explored the orthogonality of our method by combining FusedKV and FusedKV-Lite with Kivi (liu2024kivi). The experiments were conducted on a 1.5B-parameter model, applying 4-bit and 2-bit quantization to the KV cache.
@@
Model pruning reduces the number of parameters by removing weights deemed unimportant, which directly translates to lower model storage and potentially faster computation. When combining our methods with Wanda (sunsimple), we specifically pruned 50% of the Wanda model’s parameters using an unstructured approach.
@@ A.7 Long Context Performance
We evaluate models on the RULER benchmark (hsieh2024ruler). The results, presented in Table 16, demonstrate the effectiveness of our approach in preserving long-range dependencies. This stark difference highlights the effectiveness of FusedKV and FusedKV-Lite. While some information loss is inevitable with any form of cache compression, our methods demonstrate a significantly more graceful trade-off, retaining critical long-range information that is essential for practical long-context applications.

### v2

@@ 1 Introduction
Large language models (LLMs) based on the Transformer (Vaswani et al., 2017) decoder have achieved unprecedented success across a wide range of tasks and applications. However, practical deployment for real-world applications requires LLM inference to achieve low latency and high throughput (Hu et al., 2024; Ge et al., 2024; Dao et al., 2022). This becomes particularly challenging with increasingly long contexts, primarily due to the linear memory overhead of the key-value (KV) caches in autoregressive generation (Liu et al., 2024b; Adnan et al., 2024; Li et al., 2024).
To mitigate this, various methods have been proposed to reduce KV caches. Within the layer, Group-Query Attention (GQA  [^2] Unless otherwise noted, GQA mentioned in this paper uses a configuration where two query heads share a single key-value head.(Ainslie et al., 2023) and Multi-Query Attention (MQA) (Shazeer, 2019) reduce redundancy by sharing one head’s key and value across multiple query heads. MLA (Liu et al., 2024a) uses low-rank joint compression for key-value caches. Across the layer, CLA (Brandon et al., 2024) shares the KV cache between adjacent layers, while YOCO  [^3] Note that in the original YOCO paper, the cache-storing layers use efficient attention mechanisms such as Sliding Window Attention(SWA). In this paper, unless otherwise stated, we use full attention. (Sun et al., 2024) reuses the intermediate layer cache for the top-half layers. While cross-layer sharing techniques also reduce memory overhead, they consistently underperform within-layer methods, which motivates the need for a more effective cross-layer sharing strategy.
@@ 2.2 Reconstruction Function $\mathcal{F}$
the reconstruction function $\mathcal{F}_{i}$ acts as a selector, the source layer mapping $\Phi(i)$ typically contains a single index, i.e., $|\Phi(i)|=1$. Recent methods such as YOCO (Sun et al., 2024) and CLA (Brandon et al., 2024) adopt this approach.
@@ 2.4 RoPE Compatibility
To ensure that learnable weights do not disrupt the relative positional encoding property of RoPE (Su et al., 2024), we analyze the attention score. For a query at position $m$ and a key at position $n$, the attention score $A_{n,j}$ in the $j$-th 2D subspace, after incorporating a learnable weight vector ${\bm{w}}_{j}=[w_{2j},w_{2j+1}]^{T}$, can be decomposed as follows (see Appendix A.4 for the full derivation):
@@
Method Prefilling FLOPs Decoding FLOPs Cache Mem. Cache I/O.
@@ 3 Experiments
We introduce three dense language models with 332M, 650M and 1.5B parameters, both following the Qwen3 architecture (Team, 2025). More configuration are detailed in Table 6. All three models share the same configuration: a vocabulary of 128,000 tokens, a context length of 8,192, and 16 attention heads. They were trained on the FineWeb-Edu dataset (Lozhkov et al., 2024). We use the AdamW optimizer (Loshchilov and Hutter, 2017) with $\beta_{1}=0.9$, $\beta_{2}=0.95$. The learning rate followed a cosine schedule, warming up for 2,000 steps to a peak of $3\times 10^{-4}$ and decaying to a minimum of $3\times 10^{-5}$. A more detailed configuration can be found in Table 4.
@@
For model evaluation and comparison, we assess perplexity on a 500M-token validation split randomly sampled from FineWeb-Edu. We also evaluate perplexity on WikiText (Merity et al., 2016). We also conduct five-shot evaluation across diverse language tasks, including MNLI (Williams et al., 2018), SCIQ (Welbl et al., 2017), LAMBADA (LAMB-Acc) (Paperno et al., 2016), Hellaswag (Zellers et al., 2019), ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018) and MMLU (Hendrycks et al., 2021). All evaluations are based on the LM Evaluation Harness (Gao et al., ).
@@ 5.1 Cross-Layer KV Cache
KV caching is a major memory bottleneck in LLM inference (Wu et al., ; Liao and Vargas, 2024). Memory footprint can be reduced by limiting KV state retention to a subset of layers. YOCO (Sun et al., 2024) reuses a global KV cache from an intermediate layer for the model’s latter half, roughly halving memory; CLA (Brandon et al., 2024) shares caches across adjacent layers. More aggressive cross-layer reuse includes Mixture of Recurrent (Bae et al., 2025), which stacks recurrent blocks (each comprising multiple transformer layers) so later blocks reuse the first block’s KV. SVFormer (Zhou et al., 2025), which does not share the key but reuses the bottom layer’s values; and LCKV (Wu and Tu, 2024), which caches only the last layer’s KV for all earlier layers, with safeguards against cyclic dependencies. However, most approaches often overlook the distinct roles of keys and values: keys drive relevance scoring, while values supply contextual content. Treating KV as a single unit for caching neglects this functional disparity. Moreover, sharing KV states indiscriminately across layers risks undermining each layer’s specific role in hierarchical processing, potentially introducing irrelevant context and hindering its unique function. These critical aspects remain underexplored.
@@ 5.2 Early-Feature Fusion
Another line of research leverages early-layer feature fusion to stabilize gradients and enrich global features( (He et al., 2016; Huang et al., 2017; Pagliardini et al., 2024; Abdullaev and Nguyen, 2025)). ResNet (He et al., 2016) introduces residual connections that add outputs from previous layers to the current layer to stabilize training. Subsequent approaches further enhance the information flow within this framework. DenseNet (Huang et al., 2017) concatenates the outputs of all previous layers, while Denseformer (Pagliardini et al., 2024) and MambaMixer (Behrouz et al., 2024) incorporate learnable weighted coefficients into the residual paths, making the residual representations more expressive. Furthermore, reusing features from earlier layers can help alleviate the over-smoothing problem (Wang et al., 2022; Shi et al., 2022), where representations converge and lose diversity as the depth of the network increases, leading to degraded performance. To mitigate this issue, NeuTRENO (Nguyen et al., 2023) introduces a skip connection that injects a fraction of the bottom-layer output into every subsequent self-attention layer to mitigate over-smoothing. Similarly, (Abdullaev and Nguyen, 2025) enriches the transformer’s representational capacity by reusing value-residual information through a twice-attention mechanism.
@@ A.2 Architectures of Different Models
we adopt a decoder-only transformer architecture akin to Qwen3 (Team, 2025), with model sizes ranging from 332M to 4B parameters. The specific architecture of models is summarized in Table 6. The hyperparameters for Pretraining is shown in Table 4.
@@ A.6.1 Compatibility with MLA
While our proposed cross-layer fusion method, FusedKV, offers a novel way to reduce KV cache size by reconstructing information, alternative approaches achieve this through fundamental architectural modifications. A prominent example is Multi-Head Latent Attention (Liu et al., 2024a) , which reduces cache dimensionality by design rather than relying on reconstruction across layers. To investigate whether these two architextures are synergistic, we evaluate the compatibility and combined benefits of applying our method on top of an MLA-based model.
@@
A direct empirical measurement of inference speed is heavily dependent on highly optimized, hardware-specific kernels. To provide a principled analysis without the substantial engineering effort required for integration into production-grade frameworks like vLLM or SGLang, we estimated the inference speed using the open-source InferSim toolkit Team and Lab. (2025). We analyze two key metrics: Time to First Token (TTFT) for the prefill phase and Time Per Output Token (TPOT) for the decoding phase.
@@
Input/Output Length TTFT (ms) (MLA / MLA+FusedKV / MHA+FusedKV) TPOT (ms) (MLA / MLA+FusedKV / MHA+FusedKV)
@@
As established in recent work Wang et al. (2025); Chang et al. (2024), when a model’s Arithmetic Intensity(AI) exceeds the GPU’s compute-to-memory ratio, the latency from I/O communication can be effectively hidden by computation. For MLA, the AI is approximately $4\times\text{heads}$ Wang et al. (2025). To demonstrate this principle, we configured a model with 48 attention heads, creating a compute-bound scenario where the AI of MLA+FusedKV surpasses NVIDIA H20 GPU’s compute-to-memory ratio. As shown in Table 10, in this setting, the additional I/O latency from our fusion step is fully overlapped by the increased computation. Consequently, while the prefill (TTFT) benefits remain, the decoding speed (TPOT) becomes identical to that of the standard MLA model.
@@ A.6.2 Compatibility with GQA
Beyond MLA, another widely adopted architectural optimization for improving inference efficiency is Grouped-Query Attention (GQA) Ainslie et al. (2023). To investigate synergistic effects with GQA, we conduct experiments on a 650M-parameter model, combining FusedKV with a GQA architecture. All training parameters were aligned with those used in Table 2. Our results, presented in Table 11, demonstrate that FusedKV is highly compatible with GQA and their combination yields a compelling performance-cost trade-off.
@@ A.6.4 Compatibility with sliding window
A prominent hybrid architecture is the integration of Sliding Window Attention (SWA) Jiang et al. (2023) with full attention. In this section, we investigate the adaptability of FusedKV and FusedKV-Lite to a hybrid model that combines both SWA and full attention.
@@ A.6.5 COMPATIBILITY WITH OTHER EFFICIENCY TECHNIQUES
A key measure of a new efficiency technique’s utility is its ability to compose with existing methods. To evaluate the broader applicability of FusedKV and FusedKV-Lite, we investigate its compatibility with two prevalent model compression strategies: quantization (Liu et al., 2024c) and pruning (Sun et al., ).
@@
Quantization reduces memory footprint by representing weights and activations with lower-precision data types. We explored the orthogonality of our method by combining FusedKV and FusedKV-Lite with Kivi (Liu et al., 2024c). The experiments were conducted on a 1.5B-parameter model, applying 4-bit and 2-bit quantization to the KV cache.
@@
Model pruning reduces the number of parameters by removing weights deemed unimportant, which directly translates to lower model storage and potentially faster computation. When combining our methods with Wanda (Sun et al., ), we specifically pruned 50% of the Wanda model’s parameters using an unstructured approach.
@@ A.7 Long Context Performance
We evaluate models on the RULER benchmark (Hsieh et al., 2024). The results, presented in Table 16, demonstrate the effectiveness of our approach in preserving long-range dependencies. This stark difference highlights the effectiveness of FusedKV and FusedKV-Lite. While some information loss is inevitable with any form of cache compression, our methods demonstrate a significantly more graceful trade-off, retaining critical long-range information that is essential for practical long-context applications.